OpenStudy (james1107):

What is the length of the hypotenuse of a right triangle whose legs have lengths of 5 and 12? 13 15 17

OpenStudy (janu16):

Use the pythagorean theorem: a^2 + b^2 = c^2 a and b are the sides of the triangle; c is the hypotenuse. 12^2 + 5^2 = c^2 144 + 25 = c^2 169 = c^2 c = 13

OpenStudy (janu16):

do you understaand?

OpenStudy (james1107):

dang your good medal @Keigh2015 plz

OpenStudy (janu16):

do you have more?

OpenStudy (james1107):

ya

OpenStudy (janu16):

ok tag me

OpenStudy (james1107):

no need you find it as soon as i post it

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (janu16):

lol